S. Let xn3sin(0.4nT),n =0,1,2... ,24 a) What is the DFT ofx[n]? b) Plot the magnitude and phase spectra ofx[n]
Answer #1

Answer Concept DFT K-ku)-Kk+N AN 21 Asi n Couyn7)n-02, here (O)- 3Sin Q4 31 =2 N 25 25 aKo - 004 25 OP4x25 Ko 0.2xS = 5.0 k)-37.5 X=375k-s)-ks5 This is x) DFT nitude plot () Magnitude spectum 375 37.5l0ti-at 2 here a-o, b -33 5 X)=No+395) x (9= 39 Sphase spectrun /xk) in degrees Lati)= tan) here a-o, b--37 S Lx(k)= tanf 39 S.t71- tan -tan(oa) Lix t= D 90